OpenStudy (mathisfun13):

I know how to solve the problem I'm just not sure how to get q in this probability problem 42% of students at a college are full time. If 7 students are randomly selected what is the probability 4 are full-time?

OpenStudy (mathisfun13):

I know that n=7 x=4 p=.42

OpenStudy (anonymous):

p+q=1

OpenStudy (mathisfun13):

In my notes I have q=P(failure) =1-p

OpenStudy (anonymous):

yes, solve the equation i wrote for p and you'll get exactly that.

OpenStudy (mathisfun13):

so .42 plus

OpenStudy (anonymous):

oops, i mean q

OpenStudy (anonymous):

no... think is you bought something for 42 cents and you paid with a 1 dollar bill. what would your change be?

OpenStudy (mathisfun13):

.58

OpenStudy (anonymous):

that's q

OpenStudy (mathisfun13):

is that how you always get q though?

OpenStudy (anonymous):

\[P(x=4)=\left(\begin{matrix}7 \\ 4\end{matrix}\right)(.42)^4(.58)^3\]

OpenStudy (anonymous):

yep. p+q=1 so q = 1-p and p = 1-q p and q are complements in the sense that their sum is 1. this is always true for binomial probabilities.

OpenStudy (mathisfun13):

ooh okayy got it. Thankss(:

OpenStudy (anonymous):

just like n = x + (n-x)

OpenStudy (anonymous):

\[P(X=k)=\left(\begin{matrix}n \\ k\end{matrix}\right)p^kq^{n-k}\]

OpenStudy (anonymous):

you're welcome!